OpenStudy (azureilai):

How can one multiply (x^2+6)(1-x)^1/2 (basically how do you multiply quadratic equation with square root equation?)

OpenStudy (e.mccormick):

\( (x^2+6)=\sqrt{(x^2+6)^2}\)

OpenStudy (azureilai):

Can you explain how you got that please? I'm a bit confused

OpenStudy (e.mccormick):

When you put something under a root, what happens? \(2=\sqrt{4}=\sqrt{2^2}\)

OpenStudy (e.mccormick):

After getting it under a root, the rules for root multiplication can be applied.

OpenStudy (azureilai):

Ok, I see what you mean. Thanks